The relationship between the velocity,V of the wave set up in a straight string by plucking it and it's mass M, is
1
Expert's answer
2021-10-03T13:23:35-0400

The velocity of the wave in a string is given as follows (see https://courses.lumenlearning.com/suny-osuniversityphysics/chapter/16-3-wave-speed-on-a-stretched-string/):


"v = \sqrt{\dfrac{F_T}{\mu}}"

where "F_T" is the tension in the string, and "\mu = M/L" is the linear density of the string ("M" - mass of the string, "L" is its length).


Answer. "v = \sqrt{\dfrac{F_TL}{M}}".